# HiBO: Hierarchical Bayesian Optimization via Adaptive Search Space Partitioning

### v4

@@ 1 Introduction
Bayesian Optimization (BO) has emerged as a powerful approach for optimizing black-box functions that exhibit little information of its form and requires iterative evaluations for reaching the optima with significant computational cost. BO has been applied to a wide range of real-world scenarios requiring optimization over black-box functions within a limited budget of sampling times, such as Neural Architecture Search (Kandasamy et al., 2018; Ru et al., 2020), hyperparameter tuning (Hvarfner et al., 2022; Zimmer et al., 2021) and automatic database management system (DBMS) configuration tuning (Van Aken et al., 2017; Zhang et al., 2021, 2022). Typically, BO iteratively builds a probabilistic surrogate model on the objective function and uses an acquisition function to decide where to sample next, balancing exploration and exploitation to efficiently locate the optimum. However, it is well known that BO struggles to scale in high-dimensional search spaces (Binois and Wycoff, 2022; Malu et al., 2021). As dimensionality increases, the size of the search space grows exponentially and uniformly sampled points are more likely to be distant from each other (Rashidi et al., 2024). Such sparsity of nearby data increases uncertainty in the posterior distribution, which can cause the acquisition function to favor unexplored regions excessively, leading to over-exploration.
@@
Various approaches have been proposed towards mitigating the curse-of-dimensionality. Some rely on additive models applied to separate groups of dimensions (Duvenaud et al., 2011; Kandasamy et al., 2015), while others project the high-dimensional input space into a lower-dimensional subspace using linear transformations (Wang et al., 2016; Letham et al., 2020; Nayebi et al., 2019). But these methods assume specific structural properties, such as decomposability of the objective function or the existence of a hidden low-dimensional active subspace. Another line of work avoids these assumptions by restricting the sampling scope into a promising region at each iteration. Among them, some define this region as the trust region (TR) maintained around the best samples found so far (Eriksson et al., 2019; Daulton et al., 2022) while others partition the search space and focus on the most promising regions (Wang et al., 2020; Munos, 2011; Bubeck et al., 2011; Kim et al., 2020). However, these methods greedily confine the sampling scope to the only region with highest immediate potential, underutilizing the structural information generated during the partitioning process, such as the way the search space is partitioned. Such inefficient utilization of generated structural information and inflexible sampling scope confinement can lead to suboptimal performance. Moreover, with a limited initial sample budget, the constructed tree may lack reliability, and strict confinement can further prevent effective exploration of the search space. Our experiments empirically demonstrate the suboptimal performance of such approaches, where our proposed method offers $\sim$28performance improvement in real-world systems than them. While incorporating such structural information into the original BO algorithm for optimal performance introduces challenges, including designing an effective hierarchy to combine the extra module extracting structural information with the original BO, and ensuring that the structural information are utilized optimally. On the other hand, the way of partitioning the search space is critical in terms of exploitation-and-exploration trade-offs and the computational cost. Splitting the search space into coarser or fine partitions directly impacts the generated structural information and the computational overhead in this process. An adaptive mechanism for adjusting the partitioning process is needed.
@@ 2.1 Preliminaries
Bayesian Optimization (BO) is an effective technique for optimization that iteratively refines a probabilistic surrogate model of the objective function and employs an acquisition function to determine the next sampling points (Candelieri, 2021).
The goal is to solve the following problem:
$x^{*}=\underset{x\in X}{arg\>max}\;f(x)$ (1)
where $f(x)$ is a black-box function mapping the input data into the performance metric of interest. This function is typically computationally expensive to evaluate and provides limited explicit information, hence the term "black-box."
A typical BO-based algorithms begins with an initial random sampling of data points. In each iteration of its optimization loop, it fits a surrogate model to estimate the posterior probability distribution over the objective function based on the collected data. It then selects the next point maximizing the acquisition function, evaluates this point by the objective function and updates the history dataset. As iterations progress, the surrogate model becomes increasingly accurate at estimating the distribution of the objective function, leading to better sampling decisions.
@@
2.2 Related Work
Several paradigms of approaches have been developed to scale Bayesian Optimization (BO) to high-dimensional search spaces. As discussed in Section 1, some methods leverage structural assumptions about the objective function, such as additive kernels (Duvenaud et al., 2011; Kandasamy et al., 2015; Wang et al., 2018) or subspace embeddings (Wang et al., 2016; Letham et al., 2020; Nayebi et al., 2019; Papenmeier et al., 2022). Alternatively, variable-selection-based strategies are employed to identify the most important dimensions (Li et al., 2017; Spagnol et al., 2019; Shen and Kingsford, 2023; Song et al., 2022). However, these methods often rely on assumptions of the input space that may not always hold in practice, or involve uncertain user-estimated parameters for these assumptions, such as the existence of the latent subspace and its dimensionality respectively.
Another line of work focuses on optimizing the sampling scope without such assumptions, which can be further divided into 1) local modeling (Eriksson et al., 2019; Daulton et al., 2022) and 2) space partitioning (Munos, 2011; Bubeck et al., 2011; Wang et al., 2021, 2020). TuRBO, a representative of local modeling, maintains trust regions around the best samples to mitigate over-exploration. While prior space-partitioning-based methods focus on splitting the search space into different partitions and limiting the sampling scope to the most promising one. A detailed comparison space-partitioning-based optimization is presented in Table 1, which includes HOO (Bubeck et al., 2011), LA-NAS (Wang et al., 2021), LA-MCTS (Wang et al., 2020) and HiBO. HOO traditionally applies axis-aligned K-nary partitioning for splitting the space while LA-NAS extends it by leveraging linear regression to learn the linear decision boundary for bifurcation. As a more recent work, LA-MCTS applies MCTS-inspired ideas to partition the search space based on clustering and classification for more flexible space bifurcation, and becomes the state-of-the-art (SOTA) approach among space-partitioning-based work. In our work, HiBO extends these MCTS-based concepts in LA-NAS and LA-MCTS by formalizing the notion of sampling potential and employing a general search tree instead of MCTS. HiBO’s partitioning search tree can be adaptively adjusted during the search to balance exploration-exploitation trade-off, and reduce unnecessary computational cost. Additionally, HiBO does not restrict sampling scope to the only specific region, allowing greater flexibility and more comprehensive consideration collected samples.
@@ 2.2 Related Work
Table 1: Comparison between optimization algorithms based on space partitioning, where LR refers to Linear Regression and acqf. refers to acquisition function. Partitioning Op. is the operation for splitting the search space (see Section 4.1 for more details).
3 System Overview
@@
Before the optimization loop starts, similar to typical BO-based algorithms, HiBO collects an initial set of data points (history dataset) based on random sampling, ensuring the basic coverage of the search space. The initial dataset allows the global-level navigator to construct the initial search tree and enables the local optimizer to train the initial surrogate model. After the initial sampling, HiBO enters the main optimization loop. During each iteration of the loop, the global-level navigator partitions the search space based on the history dataset, and the local optimizer integrates the global-level information into its acquisition strategy for biasing the sample selection towards data points from most promising regions. The high-level workflow for each iteration is outlined below, where Figure 1 illustrates this procedure:
  • Step 1: Based on the history dataset, the global-level navigator constructs the search tree to partition the search space and generates a distribution of partition scores. While the local optimizer trains its surrogate model. The details of search space partitioning is explained in Section 4.1.
@@
The key components of HiBO mentioned in the procedure above will be explained in the following sections.
4 Global-level Navigator: Adaptive Search Space Partitioning
This section presents the design of our global-level navigator, which recursively and adaptively partitions the search space with machine learning techniques and evaluates the sampling potential of partitions balancing exploration and exploitation. We intuitively describe the sampling potential of a search space partition as the potential of sampling high-performing data points within this partition in the following sections.
@@ 4 Global-level Navigator: Adaptive Search Space Partitioning
The core motivation of introducing this global-level navigator is to partition the search space in a way maximizing the difference in sampling potential across resulting partitions. It generates a distribution of sampling potential over these partitions, which will be utilized by the local optimizer for augmenting its acquisition strategy to boost its optimization efficiency in high-dimensional search spaces (see Section 5). Besides the space partitioning itself, the design of the global-level navigator includes mechanisms for reducing computational cost and balancing trade-offs between exploration and exploitation.
4.1 Data-Driven Search-Tree-based Space Partitioning
@@
The navigator constructs a search tree in a data-driven manner, where each node represents a partition of the original search space and is associated with the samples falling into this partition. Starting from the root node representing the entire search space, this method recursively splits the space into smaller and more specific partitions, as illustrated in Step 1 of Figure 1.
@@
To partition a node, clustering techniques like K-Means (MacQueen et al., 1967) are used to group the samples from the history dataset that fall within this node’s partition into two clusters, assigning each sample a binary label. Given each sample has its cluster label, a classification model is trained on the sample features to predict the cluster labels. This classifier hence learns a decision boundary dividing the current search space partition into two parts. Each of these two parts thus forms a bifurcation of the original partition.
@@
By bifurcating the original partition into two sub-partitions, two child nodes are generated, each corresponding to one of the sub-partitions and containing the samples that fall within them. Recursively repeating this ML-based bifurcation starting from the root node will finally result in a binary tree of nodes, where each node represents a progressively refined region of the search space.
A key aspect of this approach is that clustering is performed on both sample features and their performance data. As a result, samples in the two resulting sub-partitions are expected to differ significantly in terms of both performance and feature similarity, enabling the differentiation of valuable data points within different partitions. Once the tree is fully constructed, each leaf node corresponds to a sequence of split operations, and represents a highly specific partition of the original search space. These refined partitions provide global-level insights into which regions of the search space are likely to be valuable for further sampling.
4.2 Sampling Potential Evaluation via UCT
(a) Calculation of partition scores from an example search tree.
(b) Calculation of the weighted acquisition function value corresponding to a sample point (see Section 5).
@@ 4.2 Sampling Potential Evaluation via UCT
4.3 Dynamic and Adaptive Search Tree Construction
@@
The two concerns necessitate a dynamic and adaptive strategy to control the maximum depth of the constructed search tree during optimization. Towards this target, inspired by TuRBO’s approach (Eriksson et al., 2019) for exploration control via dynamic trust region adjustment, HiBO allows the tree to expand or shrink based on consecutive successes or failures. Consecutive successes reduces the number of bifurcation and hence increases the range covered by each leaf node, thereby promoting more exploration by considering a broader set of samples with equal partition-score weights.
@@
Rule 1 (Breadth-First Search Tree Construction) Strictly follow the breadth-first order to build the tree layer by layer. Track the depth of each newly split child node and stop splitting once they reaches the current maximum tree depth, which is shown by the red line in Figure 2(a).
Rule 2 (Adaptive Maximum Tree Depth) After a certain number of consecutive successes (improvements over the current best), reduce the maximum tree depth to broaden the range covered by each leaf node. Conversely, increase the maximum depth after consecutive failures for more focused biasing and hence more exploitation.
@@
5 Local Optimizer: Partition-Score-Weighted Acquisition
@@
Given all leaf nodes represent the entire search space, each leaf node serves as an information source of the global partitioning. The UCT score of each leaf node reflects the sampling potential of each partition of the whole search space and should be effectively utilized. As illustrated in Figure 2(a), after the tree is built, each leaf node $j$ is assigned with a UCT score $UCT_{j}$ following Equation 2. To scale all scores into a positive and reasonable range, we calculate the partition score $P_{i}$ for each node $\Omega_{i}$ by a temperature-controlled softmax function:
@@
Given the partition score, the local optimizer modifies its the acquisition function value $\alpha(x;D)$ for $i$-th candidate data point $x_{i}$, conditioned on the history dataset $D$, as shown in Figure 2(b) and below:
$\alpha_{P}(x_{i};D)=\alpha(x_{i};D)\cdot P_{leaf_{x_{i}}}$ (4)
where $leaf_{\bf{x}}$ denotes the leaf node containing $\bf{x}$. With the weighted acquisition function value, the next sample $x_{next}$ to be evaluated is the one maximizing it:
$x_{next}=\underset{x\in X}{arg\>max}\;\alpha_{P}(x;D)$ (5)

### v6

# Navigating in High-Dimensional Search Space: A Hierarchical Bayesian Optimization Approach

@@ 1 Introduction
Bayesian Optimization (BO) is a powerful method for optimizing black-box functions whose internal structure is largely unknown and requires iterative evaluations for reaching the optima with significant computational cost. It has seen broad adoption in real-world scenarios, including Neural Architecture Search (Kandasamy et al., 2018; Ru et al., 2020), hyperparameter tuning (Hvarfner et al., 2022; Zimmer et al., 2021), and automatic configuration of database management systems (Van Aken et al., 2017; Zhang et al., 2021, 2022). However, it is well known that BO struggles to scale in high-dimensional search spaces due to reasons including the exponentially increased search space and over-exploration (Binois and Wycoff, 2022; Malu et al., 2021; Rashidi et al., 2024).
@@
Various methods mitigate the curse of dimensionality by restricting sampling to the most promising region (Eriksson et al., 2019; Daulton et al., 2022; Wang et al., 2020; Munos, 2011; Bubeck et al., 2011; Kim et al., 2020). However, such greedy confinement can lead to inefficient use of structural information (e.g., how partitions are formed) and inflexible sampling scope, resulting in suboptimal performance—particularly when initial budgets are small. Fully exploiting these structural insights necessitates a hierarchical approach that integrates partitioning modules, enabling a balanced exploration-exploitation trade-off within feasible computational overheads. At the same time, the granularity of the partition—whether coarser or finer—directly affects both the quality of the extracted information and the overhead of maintaining it. This interplay underscores the need for an adaptive partitioning mechanism: coarse partitions may overlook fine-grained optima, whereas overly fine ones risk high costs and premature overfitting to limited data.
@@ 2.1 Preliminaries
Bayesian Optimization (BO) iteratively refines a probabilistic surrogate of a black-box objective function and selects new samples via an acquisition function (Candelieri, 2021). Formally, it aims to solve:
$x^{*}=\arg\max_{x\in X}\,f(x),$ (1)
where $f(x)$ is expensive to evaluate and reveals little explicit structure. BO typically starts with a few random points, fits a surrogate model to the observed data, and then chooses the next query by maximizing an acquisition function. As more data is gathered, the surrogate improves, guiding increasingly effective sampling decisions.
@@
Table 1: Comparison between optimization algorithms based on space partitioning, where LR refers to Linear Regression and acqf. refers to acquisition function. Partitioning Op. is the operation for splitting the search space (see Section 3.2.1 for more details).
2.2 Related Work
Several paradigms of approaches have been developed to scale Bayesian Optimization (BO) to high-dimensional search spaces. Some methods leverage structural assumptions about the objective function, such as additive kernels (Duvenaud et al., 2011; Kandasamy et al., 2015; Wang et al., 2018) or subspace embeddings (Wang et al., 2016; Letham et al., 2020; Nayebi et al., 2019; Papenmeier et al., 2022). Alternatively, variable-selection-based strategies are employed to identify the most important dimensions (Li et al., 2017; Spagnol et al., 2019; Shen and Kingsford, 2023; Song et al., 2022). However, these assumptions of the input space may not always hold in practice, or involve uncertain user-estimated parameters for these assumptions, such as the existence of the latent subspace and its dimensionality respectively.
Another line of work focuses on optimizing the sampling scope without such assumptions, which can be further divided into 1) local modeling (Eriksson et al., 2019; Daulton et al., 2022) and 2) space partitioning (Munos, 2011; Bubeck et al., 2011; Wang et al., 2021, 2020). TuRBO, a representative of local modeling, maintains trust regions around the best samples to mitigate over-exploration. While prior space-partitioning-based methods focus on splitting the search space into different partitions and limiting the sampling scope to the most promising one. A detailed comparison space-partitioning-based optimization is presented in Table 1, which includes HOO (Bubeck et al., 2011), LA-NAS (Wang et al., 2021), LA-MCTS (Wang et al., 2020) and HiBO. HOO traditionally applies axis-aligned K-nary partitioning for splitting the space. As more recent work, LA-NAS and LA-MCTS apply MCTS-inspired ideas to partition the search space based on machine learning techniques for more flexible space bifurcation. In our work, HiBO extends these MCTS-based concepts in LA-NAS and LA-MCTS by formalizing the notion of sampling potential and employing a general search tree instead of MCTS. HiBO’s partitioning search tree can be adaptively adjusted during the search to balance exploration-exploitation trade-off, and reduce unnecessary computational cost. Additionally, HiBO does not restrict sampling scope to the only specific region, allowing greater flexibility and more comprehensive consideration collected samples.
3 Methodology: Hierarchical Bayesian Optimization
3.1 Overview
@@
After the initial sampling, HiBO enters the main optimization loop. The high-level workflow for each iteration is outlined below, where Figure 1 illustrates this procedure:
  • Step 1: Based on the history dataset, the global-level navigator constructs the search tree to partition the search space and generates a distribution of partition scores, while the local optimizer trains its surrogate model. The details of search space partitioning is explained in Section 3.2.1.
@@
The key components of HiBO mentioned in the procedure above will be explained in the following subsections.
3.2 Global-level Navigator: Adaptive Search Space Partitioning
@@
The core motivation of introducing this global-level navigator is to partition the search space in a way maximizing the difference in sampling potential  [^1] Note that we intuitively describe the sampling potential of a search space partition as the potential of sampling high-performing data points within this partition in the following sections. across resulting partitions. The navigator generates a distribution of sampling potential over these partitions, which will be utilized by the local optimizer for augmenting its acquisition strategy to boost its optimization efficiency in high-dimensional search spaces (see Section 3.3). Besides the space partitioning itself, the design of the global-level navigator includes mechanisms for reducing computational cost and balancing trade-offs between exploration and exploitation.
3.2.1 Data-Driven Search-Tree-based Space Partitioning
@@
During each iteration, the navigator constructs a search tree in a data-driven manner, where each node represents a partition of the original search space and is associated with the samples falling into this partition. Starting from the root node representing the entire search space, this method recursively splits the space into smaller and more specific partitions, as illustrated in Step 1 of Figure 1.
@@
To partition a node, clustering techniques like K-Means (MacQueen et al., 1967) are used to group the samples from the history dataset that fall within this node’s partition into two clusters, assigning each sample a binary label. A classification model is then trained on sample features to predict the cluster labels and hence learns a decision boundary dividing the current search space partition into two parts. Each of these two parts thus forms a bifurcation of the original partition. Considering clustering is performed on both sample features and their performance data, samples in the two resulting sub-partitions are expected to differ significantly in terms of both performance and feature similarity, enabling the differentiation of valuable data points within different partitions.
@@
By bifurcating the original partition into two sub-partitions, two child nodes are generated, each corresponding to one of the sub-partitions and containing the samples that fall within them. Recursively repeating this ML-based bifurcation starting from the root node will finally result in a binary tree of nodes, where each leaf node corresponds to a sequence of split operations and represents a highly specific partition of the original search space. These refined partitions provide global-level insights into which regions of the search space are likely to be valuable for further sampling.
3.2.2 Sampling Potential Evaluation via UCT
@@
3.2.3 Dynamic and Adaptive Search Tree Construction
@@
The two concerns necessitate a dynamic and adaptive strategy to control the maximum depth of the constructed search tree during optimization. Towards this target, inspired by TuRBO’s approach (Eriksson et al., 2019) for exploration control via dynamic trust region adjustment, HiBO allows the tree to expand or shrink based on consecutive successes or failures. Consecutive successes reduce the number of bifurcation and hence increases the range covered by each leaf node, thereby promoting more exploration by considering a broader set of samples with equal partition-score weights.
@@
  • Rule 1 (Breadth-First Search Tree Construction) Strictly follow the breadth-first order to build the tree layer by layer. Track the depth of each newly split child node and stop splitting once they reaches the current maximum tree depth, which is shown by the red line in Figure 2(a);
  • Rule 2 (Adaptive Maximum Tree Depth) After a certain number of consecutive successes (improvements over the current best), reduce the maximum tree depth to broaden the range covered by each leaf node. Conversely, increase the maximum depth after consecutive failures for more focused biasing and hence more exploitation;
@@
(a) Calculation of partition scores from an example search tree.
(b) Calculation of the weighted acquisition function value corresponding to a sample point (see Section 3.3).
@@
3.3 Local Optimizer: Partition-Score-Weighted Acquisition
@@
After the tree is built, all leaf nodes together represent the entire search space, with each of them serving as an information source of the global partitioning. Each leaf node $j$ is assigned with a UCT score $UCT_{j}$ following Equation 2, reflecting the sampling potential of each partition of the whole search space. For effectively utilizing them, we calculate the partition score $P_{i}$ for each node $\Omega_{i}$ by a temperature-controlled softmax function to scale all scores into a positive and reasonable range:
@@
Given the partition score, the local optimizer modifies its the acquisition function value $\alpha(x;D)$ for $i$-th candidate data point $x_{i}$, conditioned on the history dataset $D$, as shown in Figure 2(b):
$\alpha_{P}(x_{i};D)=\alpha(x_{i};D)\cdot P_{leaf_{x_{i}}}$
$leaf_{\bf{x}}$ denotes the leaf node containing $\bf{x}$. With the weighted acquisition function value, the next sample $x_{next}$ to be evaluated is the one maximizing:
$x_{next}=\underset{x\in X}{arg\>max}\;\alpha_{P}(x;D)$
@@
4 Theoretical Analysis
This section analyzes HiBO’s theoretical properties, focusing on sample efficiency and scalability. We derive an upper bound for samples in the most promising partition and discuss computational complexity. More discussions on comparisons with prior works (Wang et al., 2021, 2020) and scalability are provided in Appendix B.
4.1 Bounding the Most Promising Partition
At the root node of the partition tree (the global search domain $\Omega$), we define a partition error $\delta$ to capture how far the split deviates from a perfect $N/2$$N/2$ division when separating samples with higher and lower performance.
Formally, $\delta$ is the number of samples assigned to the subset with lower average performance that actually belong to the higher-performance set (or vice versa). By Assumption (4) in Appendix B.1, it can be proved that at most $\delta$ samples fall into such a misclassified subpartition, so that each partition is of size at most $\frac{N}{2}\;+\;\delta$. We now extend this reasoning to a tree of height $h$, constructed by recursively subdividing the most promising region at each node based on the global-level navigator’s UCT-driven scores:
Theorem 4.1.
(Bound on Most Promising Leaf). Consider a partition tree of height $h$ built by recursively splitting the sub-domain with the highest partition score. Let $\delta_{\max}$ be the maximum partition error of any node along that “leftmost” path. Then the sub-domain (leaf) at depth $h$ on that path contains at most $2\,\delta_{\max}\,\Bigl{(}1-\frac{1}{2^{h}}\Bigr{)}\;+\;\frac{N}{2^{h}}$ solutions.
Proof.
[…]
$\underbrace{\delta_{h}\;+\;\frac{\delta_{h-1}}{2}\;+\;\frac{\delta_{h-2}}{2^{2}}\;+\;\dots\;+\;\frac{\delta_{1}}{2^{h-1}}}_{\text{accumulated partition error}}\;+\;\frac{N}{2^{h}}.$
Defining $\delta_{\max}=\max\{\delta_{1},\delta_{2},\dots,\delta_{h}\}$ and bounding each $\delta_{j}$ by $\delta_{\max}$ yields:
$\delta_{h}+\frac{\delta_{h-1}}{2}+\dots+\frac{\delta_{1}}{2^{h-1}}$ $\;\;\leq\;\;\delta_{\max}\,\Bigl{(}1\;\;+\dots\;\frac{1}{2^{\,h-1}}\Bigr{)}$
= $\;=\;2\,\delta_{\max}\,\Bigl{(}1-\frac{1}{2^{h}}\Bigr{)}.$
Combining this with $\tfrac{N}{2^{h}}$ completes the proof,
[…]
The bound above shows that, as $h$ increases, the size of the subpartition containing the most promising solutions shrinks at a rate $\tfrac{N}{2^{h}}$, up to an additive term that depends on the worst-case partition error $\delta_{\max}$. Hence, in an idealized setting with small $\delta_{\max}$, HiBO zooms in on a small fraction of $\Omega$ at exponential speed in terms of $h$. The assumption can be found in Appendix B.1.
4.2 Implications
Exponential Pruning of Sub-Optimal Regions. Theorem 4.1 shows HiBO can isolate high-performing samples at a rate of $\tfrac{N}{2^{h}}$ with tree depth $h$. While computational cost grows exponentially with $h$, the quality of the best partition approaches the true optimum in $\Omega$, accelerating pruning compared to random sampling’s $\frac{N}{2}$ complexity.
Impact of Accurate Partitioning ($\delta_{\max}$). The partition accuracy $\delta_{\max}$ improves as more observations refine the UCT scores. This explains HiBO’s enhanced performance in larger search spaces: once reliable UCT estimates are established, the advantage over uniform exploration becomes more significant.